OpenStudy (anonymous):

How can 1/2x − 5 = 1/3x + 6 be set up as a system of equations

jigglypuff314 (jigglypuff314):

it's the y = y idea for example if given x+3 = 2x+7 then y = x+3 and y = 2x+7

OpenStudy (anonymous):

^^^^^

OpenStudy (anonymous):

how is that supposed to help

OpenStudy (jdoe0001):

\(\bf {\color{red}{ y}}=\cfrac{1}{2}x-5\qquad \qquad {\color{red}{ y}}=\cfrac{1}{3}x+6 \\ \quad \\ {\color{red}{ y}}={\color{red}{ y}}\implies ?\)

OpenStudy (anonymous):

how Is that suppose to mean

jigglypuff314 (jigglypuff314):

he basically gave you what you wanted... system of equations: y = 1/2 x - 5 y = 1/3 x + 6

OpenStudy (anonymous):

the answer choices are A. 2y + x = −10 3y + x = 18 B. 2y + 2x = −10 3y + 3x = 18 C. 2y − x = −10 3y − x = 18 D. 2y − 2x = −10 3y − 3x = 18

jigglypuff314 (jigglypuff314):

y = 1/2 x - 5 multiply through by 2 y = 1/3 x + 6 multiply through by 3